# Quantum dot thermal machines – a guide to engineering

### v1

@@ I Introduction
Quantum thermal machines, heat engines and refrigerators, have been a focus of intense research in the past decades, both as a means of gleaning insight into thermodynamics in the quantum regime Campbell et al. (2025), and for practical applications. Miniaturisation has been a central theme in contemporary technological development, and compact and efficient sources of power and cooling has become an essential requirement for the development of quantum technologies.
From a practical viewpoint, out of the wide variety of theoretical configurations of quantum thermal machines Cangemi et al. (2024), and systems in which they have been realised Myers et al. (2022): from lasers and solar cells, to superconducting circuits Aamir et al. (2025), qubits, levitated nanoparticles Kuhn et al. (2017), and a single atom Roßnagel et al. (2016); for near-term applications the most promising is a continuous electronic particle-exchange heat engine Kosloff and Levy (2014). These devices are autonomous, require no feedback, no time-dependent driving, can be fabricated harnessing existing technological achievements of nanoscale electronics, and readily incorporated with other functional elements.
The simplest example of such a heat engine is a single electron transistor based on a single quantum dot. Despite its simplicity, such a device has been shown to demonstrate efficiency of over $\eta=0.7$ of the Carnot efficiency Josefsson et al. (2018), as well as agreement with the Curzon-Ahlborn efficiency Novikov (1958); Curzon and Ahlborn (1975) at maximum power. These experimental results agree with the theoretical prediction of system with infinitely narrow transition energy windows optimising thermoelectric efficiency Mahan and Sofo (1996); Whitney (2014). Generally, SET-based quantum dot heat engines have been predicted to reach thermodynamic limits Esposito et al. (2009); Josefsson et al. (2019).
More recently, a similar experiment with SET-based heat engine has been performed, employing a single molecule as a working medium Volosheniuk et al. (2025). It showed comparably high performance, but also a dependence of heat engine operation on the magnetic field – the suppression of the Kondo effect and the magnetic shift of the energy levels in the molecule reduced both the efficiency of the heat engine and the maximum output power.
These results raise a general question that, to our knowledge, has not been considered before – how do the internal dynamics of the quantum dot in an SET thermal machine affect its performance, and can it be improved by optimising them? Or, what properties should a quantum dot have to make the best-performing thermal machine?
@@
To limit our considerations to both promising and the experimentally tractable, we consider general devices in the sequential tunnelling regime with the energy level spacing much smaller than $k_{\rm{B}}T$, so that the energy transfer can be considered narrow, in line with the prediction of optimal thermoelectric performance Mahan and Sofo (1996); Whitney (2014) (this regime frequency applies to single-molecule devices Pyurbeeva et al. (2021)). We also will not consider passive heat flow or vibrational effects. The reason for the latter is that, except for highly specialised cases such as phonon-assisted tunnelling Sowa et al. (2017), both reduce performance, and are tunable independently of electronic parameters, and thus, for practical applications, can be mitigated separately.
This paper is structured as follows: in Section II, we describe all configurations of a SET and identify the regimes where it operates as a heat engine and refrigerator; in Section 3, we derive the expressions for efficiency and power of an SET heat engine and refrigerator in the linear regime; in Section 4, we consider, in order, the effect of the entropy difference between the two change states involved in transport on thermal machine operation, the same effect in the case of normalised conductance, and the effect of detailed balance breaking. Finally, before summarising in Discussion, in Section 5, we study the effects of the entropy difference, tunnel coupling asymmetry, and detailed balance breaking on the noise in the system.
To summarise, we demonstrate that for a narrow energy transmission band, the dynamics of the system can be described through four characteristic parameters that can be reduced to two if noise is not considered, and which can act as a guide to engineering quantum states for optimisation of nanoscale thermal machines.
@@ II Thermal machine configurations
where $\Delta S_{u}$ is the entropy change of the universe, $\mu_{H/C}$ and $T_{H/C}$ are the chemical potentials and temperatures of the hot/cold baths, and $\epsilon$ is the charging energy corresponding to current reversal. This gives the current reversal level energy:
@@ III.2 Refrigerator
In the linear regime, where $T_{H}\approx T_{C}\approx T$, and $\Delta V$ and $\Delta T$ are small, this lead to operation conditions for $\varepsilon=\epsilon-\mu$: $\varepsilon\in\{0;(\Delta V/\Delta T)T\}$, and $\varepsilon$ can be comparable to $T$ is $\Delta V$ and $\Delta T$ are of similar magnitudes.
@@ IV.1 Entropy difference
If electron transport is mediated by a single narrow energy level with no additional dissipation, it can be fully described by four exchange rates between the QD and the electrodes (Fig.4a). These rates contain all the information on the internal dynamics of the quantum dot that affects transport, and depend on the densities of states of the electrodes at a single value of $\varepsilon$. We will use them to find $G(\varepsilon)$, the conductance of the QD, and with it, its efficiency and power operating as a heat engine and refrigerator, using the results from the previous section.
@@
as the characteristic of the device, where $A$ is the “amplitude” coefficient.
@@
Fig.4c shows the standard power-efficiency plots for an SET heat engine with varying values of $\Delta S$. It should be noted that, unlike the typical diagrams, where the “loop” Josefsson et al. (2018) tends to zero after reaching maximum efficiency at non-zero power, here, due to the absence of passive heat flow between the hot and cold baths in the model, maximum efficiency is reached at zero power, and the familiar “loop” shape for each non-zero value of $\Delta S$ is made up of the different magnitudes of the positive and negative sections of the thermocurrent $L$.
It can be seen that increasing $\Delta S$ leads to and increase in both power and efficiency throughout the range of gate tuning $\varepsilon$ (Fig.4c), leading to greater maximum power and maximum efficiency. Figure 4d shows maximum power plotted against efficiency at maximum power for a range of load resistances, and demonstrates that $\Delta S$ increases maximum output power at all heat engine operation regimes.
@@ IV.2 Normalised conductance
Figure 5: a) Normalised gate-dependent conductance (i) and thermoelectric susceptibility for several values of $\Delta S$. b) The power-efficiency diagrams for various values of $\Delta S$ (the positive and negative values of $\Delta S$ have the same dependence). The non-degenerate case is shown in red. c) The parameter plots of maximum power vs. efficiency at maximum power at varying load resistances for several values of $\Delta S$. d) The maximum cooling power of an SET refrigerator with normalised conductance as a function of temperature difference for several values of $\Delta S$. e) Cooling power of an SET refrigerator with normalised conductance as a function of $\Delta S$ at fixed $\Delta T$ and $\Delta V$.
@@
and the respective $L$ can provide a general description of transport properties of an arbitrary nanodevice with a narrow transmission band – such an approach had been used in Volosheniuk et al. (2025) to describe an SET heat engine where both conductance and thermocurrent were significantly affected by the Kondo effect and therefore could not be readily described by the simple rate equation approach.
@@
To explore this effect, in Fig.5e, we plot the dependence of maximum cooling power at a function of $\Delta S$ for different values of operating bias voltage. The entropy sign opposite to the applied bias dominates, meaning that for hole-dominated configuration (Fig.3b(ii)), negative values of $\Delta S$ are preferable, while for an electron-dominated configuration (Fig.3b(i)) the opposite holds true. However, the functions $Q_{max}(\Delta S)$ show distinct maxima, meaning that, unlike the heat engine case, where an increase of $\Delta S$ is always beneficial, even for normalised conductance, for an SET operating as a refrigerator, there is an optimal value of $\Delta S$ that maximises cooling power for given operating conditions.
@@ IV.3 Detailed balance breaking
Another aspect of the internal dynamics of a quantum dot that can effect its performance as a thermal machine is detailed balance breaking. Such effects, as well as an advantage in heat engine performance, have been demonstrated in spintronics Katcko et al. (2019); Cho (2022). Molecules demonstrating chirality-induced spin-selectivity effect (CISS) Naaman et al. (2019); Bloom et al. (2024) also comprise a promising platform for realising novel nanoscale thermal machines breaking detailed balance.
@@
The symmetry breaking between the positive and negative current directions means that charge will be accumulated across the SET, resulting in an induced offset bias voltage in equilibrium (at zero current).
@@
This means that tuning the detailed balance breaking and tunnel coupling asymmetry coefficients to increase $\xi$ improves the device’s performance as a heat engine. Its performance as a refrigerator can also be optimised, but the optimal value of $\sigma$ depends of the operating voltage and temperature difference. Finally, despite the arguments for the analysis with normalised conductance, it should also be noted that increasing $\xi$ increases overall conductance and therefore thermal machine performance.
@@ V Noise and constancy
The final significant characteristic of a thermal machine that had not been considered so far is noise, or the constancy of power delivery or extraction. Unlike the power and efficiency, both based on mean current through the device, in which all three asymmetry parameters: $\Delta S$, $\gamma$, and $\alpha$, came in in a single quantity, noise, or current fluctuations, has a more direct dependence on the microscopic dynamics of the system, and the three asymmetry parameters, $\Delta S$, $\gamma$, and $\alpha$ all affect it differently.
A general expression for current fluctuations, $\chi$, as a function of four tunnel rates for a SET has the form Landi et al. (2024):
@@
where, for brevity, we define: $\Gamma_{L/R}=\Gamma_{TL/R}+\Gamma_{FL/R}$. The fluctuations of power are then equal to $\varepsilon\chi$.
However, fluctuations of current in the output circuit are detrimental for thermal machine operation even if they are associated with little power (around $\varepsilon=0$), so for our analysis, we will consider the fluctuations of current directly.
Figure 6a shows the effect of all three asymmetry parameters on the power vs. noise diagrams for a heat engine for a normalised conductance peak height.
@@
A non-zero entropy difference between the charge states, $\Delta S$ splits the dependence for a non-degenerate level into two branches – one with higher power and one with lower for the same noise levels, corresponding to the asymmetric peaks of the thermocurrent (Fig.6a(i)). The tunnel coupling asymmetry coefficient, $\gamma$, affects the power vs. noise relation is a much less profound way (Fig.6(ii)), as, in the absence of detailed balance breaking, it doesn’t affect conductance, and thus output power, but it can be observed that absolute large values of $\gamma$ reduce current fluctuations to a small extent. Finally, the introduction of detailed balance breaking, described by non-zero values of $\alpha$, leads to a dramatic increase of current fluctuations, even in the absence of concurrent tunnel coupling asymmetry and thus the same power output (Fig.6a(iii)). A non-zero value of $\gamma$ would lead to an advantage in power, similarly to an increase of entropy difference.

### v2

@@ I Introduction
Quantum thermal machines, heat engines and refrigerators, have been the focus of intense research in the past decades, both as a means of gleaning insight into thermodynamics in the quantum regime Campbell et al. (2025), and for practical applications. Miniaturisation is a central theme in contemporary technological development, and compact and efficient sources of power and cooling have become an essential requirement for the development of quantum technologies.
From a practical viewpoint, out of the wide variety of theoretical configurations of quantum thermal machines Cangemi et al. (2024), and systems in which they have been realised Myers et al. (2022); Kosloff (2024): from lasers and solar cells, to superconducting circuits Aamir et al. (2025), qubits, levitated nanoparticles Kuhn et al. (2017), and a single atom Roßnagel et al. (2016); for near-term applications the most promising is a continuous electronic particle-exchange heat engine Humphrey et al. (2002); Humphrey and Linke (2005); Kosloff and Levy (2014). These devices are autonomous, require no feedback, no time-dependent driving, can be fabricated harnessing existing technological achievements of nanoscale electronics, and readily incorporated with other functional elements.
The simplest example of such a heat engine is a single electron transistor based on a single quantum dot. Despite its simplicity, such a device has been shown to demonstrate efficiency of over $\eta=0.7$ of the Carnot efficiency Josefsson et al. (2018), as well as agreement with the Curzon-Ahlborn efficiency Novikov (1958); Curzon and Ahlborn (1975) at maximum power. These experimental results follow the theoretical prediction of a system with an infinitely narrow transmission energy bands window optimising thermoelectric efficiency Mahan and Sofo (1996); Whitney (2014). Generally, SET-based quantum dot heat engines have been predicted to reach thermodynamic limits Esposito et al. (2009); Josefsson et al. (2019).
More recently, a similar experiment with an SET-based heat engine has been performed, employing a single molecule as a working medium Volosheniuk et al. (2025). It showed comparably high performance, but also a dependence of heat engine operation on the magnetic field – the suppression of the Kondo effect and the magnetic shift of the energy levels in the molecule reduced both the efficiency of the heat engine and the maximum output power.
These results demonstrate that for a quantum dot with a more complex internal structure than a single energy level with a two-fold spin degeneracy, this internal structure, or microscopic dynamics: additional spatial degeneracy from symmetry, energy level structure, spin interactions, etc, directly affect its performance as a thermal machine.
This raises a general question which, to our knowledge, has not been considered before – how do the internal dynamics of the quantum dot in an SET thermal machine affect its performance, and can it be improved by optimising them? Or, what properties should a quantum dot have to make the best-performing thermal machine?
@@
To limit our considerations to both promising and the experimentally tractable, we consider general devices in the sequential tunnelling regime with the energy level spacing much smaller than $k_{\rm{B}}T$, so that the energy transfer window can be considered narrow (the narrow band approximation), in line with the prediction of optimal thermoelectric performance Mahan and Sofo (1996); Whitney (2014) (this regime frequently applies to single-molecule devices Pyurbeeva et al. (2021)). We also will not consider passive heat flow or vibrational effects. The reason for the latter is that, except for highly specialised cases such as phonon-assisted tunnelling Sowa et al. (2017), both reduce performance, and are tunable independently of electronic parameters, and thus, for practical applications, can be mitigated separately.
In the narrow energy band approximation, we treat the unknown internal dynamics of a quantum dot through three asymmetry parameters – the asymmetry between adding an electron to the QD and removing it from it, relating to the relative number of available quantum states, or the difference in entropy between the charge states, $\Delta S$; the asymmetry between the tunnel couplings of the QD to the baths, $\gamma$; and $\alpha$, charactering the degree of detailed balance breaking in the system. The fourth and final parameter is the overall coupling strength.
To summarise, we demonstrate that in the narrow-band approximation, the dynamics of the system can be described through four characteristic parameters, three asymmetries and an overall scaling. We analyse the performance of an SET based on a quantum dot with non-trivial internal dynamics, including in the case of detailed balance breaking, and demonstrate that power and efficiency properties can be characterised through two parameters – the coupling strength and a combination of three asymmetries, while noise, or the constancy of power delivery or extraction, depends on all four parameters independency. We believe our results can act as guidelines to engineering quantum states for optimisation of nanoscale thermal machines.
This paper is structured as follows: in Section II, we describe all configurations of a SET and identify the regimes where it operates as a heat engine and refrigerator; in Section 3, we derive the expressions for efficiency and power of an SET heat engine and refrigerator in the linear regime; in Section 4, we consider, in order, the effect of the entropy difference between the two charge states involved in transport on thermal machine operation, the same effect in the case of normalised conductance, and the effect of detailed balance breaking. Finally, before summarising in Discussion, in Section 5, we study the effects of the entropy difference, tunnel coupling asymmetry, and detailed balance breaking on the noise in the system.
@@ II Thermal machine configurations
where $\Delta S_{u}$ is the total entropy change of the universe (that of the QD and both baths), $\mu_{H/C}$ and $T_{H/C}$ are the chemical potentials and temperatures of the hot/cold baths, and $\epsilon$ is the charging energy corresponding to current reversal. This gives the current reversal level energy Humphrey et al. (2002):
@@ III.2 Refrigerator
In the linear regime, where $T_{H}\approx T_{C}\approx T$, and $\Delta V$ and $\Delta T$ are small, this leads to operation conditions for $\varepsilon=\epsilon-\mu$: $\varepsilon\in\{0;(\Delta V/\Delta T)T\}$, and $\varepsilon$ can be comparable to $T$ is $\Delta V$ and $\Delta T$ are of similar magnitudes.
@@ IV.1 Entropy difference
If electron transport is mediated by a single narrow energy band with no additional dissipation, it can be fully described by four exchange rates between the QD and the electrodes (Fig.4a). These rates contain all the information on the internal dynamics of the quantum dot that affects transport, and depend on the densities of states of the electrodes at a single value of $\varepsilon$. We will use them to find $G(\varepsilon)$, the conductance of the QD, and with it, its efficiency and power operating as a heat engine and refrigerator, using the results from the previous section.
@@
as the characteristic of the device, where $A$ is the “amplitude” coefficient, in the case the rates defined in Eq.14, $A=(d_{F}/T)\cdot\gamma_{R}\gamma_{L}/(\gamma_{R}+\gamma_{L})$ (Eq.15).
@@
Fig.4c shows the standard power-efficiency plots for an SET heat engine with varying values of $\Delta S$. It should be noted that, unlike the typical diagrams, where the “loop” Josefsson et al. (2018) tends to zero after reaching maximum efficiency at non-zero power, here, due to the absence of heat leakage, passive heat flow between the hot and cold baths in the model, maximum efficiency is reached at zero power, and the familiar “loop” shape for each non-zero value of $\Delta S$ is made up of the different magnitudes of the positive and negative sections of the thermocurrent $L$.
It can be seen that increasing $\Delta S$ leads to an increase in both power and efficiency throughout the range of gate tuning $\varepsilon$ (Fig.4c), leading to greater maximum power and maximum efficiency. Figure 4d shows maximum power plotted against efficiency at maximum power for a range of load resistances, and demonstrates that $\Delta S$ increases maximum output power at all heat engine operation regimes.
@@ IV.2 Normalised conductance
Figure 5: a) Normalised gate-dependent conductance (i) and thermoelectric susceptibility for several values of $\Delta S$. b) The power-efficiency diagrams for various values of $\Delta S$ (the positive and negative values of $\Delta S$ have the same dependence). The non-degenerate case is shown in red. c) The parameter plots of maximum power vs. efficiency at maximum power at varying load resistances for several values of $\Delta S$. d) The maximum cooling power of an SET refrigerator with normalised conductance as a function of temperature difference for several values of $\Delta S$. e) Cooling power of an SET refrigerator with normalised conductance as a function of $\Delta S$ at fixed $\Delta T$ and $\Delta V$. Similarly to Fig.4, the plots are designed to illustrate the qualitative behaviour and thus the their scale can very.
@@
where $A$ and $\sigma$ are effectively fitting parameters quantifying the height of the conductance peak and the thermocurrent asymmetry. This form of $G(\varepsilon)$ and the respective $L(\varepsilon)$ can provide a general description of transport properties of an arbitrary nanodevice with a narrow transmission band – such an approach had been used in Volosheniuk et al. (2025) to describe an SET heat engine where both conductance and thermocurrent were significantly affected by the Kondo effect and therefore could not be readily described by the simple rate equation approach. The asymmetry parameter $\sigma$ no longer has an immediate physical interpretation in this case.
@@
To explore this effect, in Fig.5e, we plot the dependence of maximum cooling power at a function of $\Delta S$ for different values of operating bias voltage. The entropy sign opposite to the applied bias dominates, meaning that for hole-dominated configuration (Fig.3b(ii)), negative values of $\Delta S$ are preferable, while for an electron-dominated configuration (Fig.3b(i)) the opposite holds true. However, the functions $Q_{max}(\Delta S)$ show distinct maxima, meaning that, unlike the heat engine case, where an increase of $\Delta S$ is always beneficial, even for normalised conductance, for an SET operating as a refrigerator, there is an optimal value of $\Delta S$ that maximises cooling power for given operating conditions. The reason for the existence of an optimum value of $\Delta S$ for a refrigerator lies in the finite range of $\varepsilon$ corresponding to the refrigeration regime of the SET (Eq.10). The maximum of the conductance peak, $G(\varepsilon)$ (Eq.16) lies at $\varepsilon=T\Delta S/2$, and thus, large values of $\Delta S$ can move the conductance peak outside the operation window, leading to a decrease in the device’s performance.
@@ IV.3 Detailed balance breaking
Another aspect of the internal dynamics of a quantum dot that can effect its performance as a thermal machine is detailed balance breaking. It has been experimentally observed in a wide variety of physical systems with broken time-reversal symmetryEntin-Wohlman et al. (1995); Nakamura et al. (2010); Zhu and Fan (2014); Khandekar et al. (2020); Katcko et al. (2019); Cho (2022), and has been proved to be compatible with thermodynamics Alicki et al. (2023); Blum and Gelbwaser-Klimovsky (2025). It has also been shown that in devices based on spintronic spin-selectivity detailed balance breaking leads to an advantage in heat engine performance Katcko et al. (2019); Cho (2022). Molecules demonstrating the recently-proposed chirality-induced spin-selectivity effect (CISS) Naaman et al. (2019); Bloom et al. (2024) also comprise a promising potential platform for realising novel nanoscale thermal machines breaking detailed balance.
@@
The symmetry breaking between the positive and negative current directions means that charge will be accumulated across the SET, resulting in an induced offset bias voltage in equilibrium (at zero current), $\Delta V$, equal to $2T\alpha$.
@@
This means that tuning the detailed balance breaking and tunnel coupling asymmetry coefficients to increase $\xi$ improves the device’s performance as a heat engine. This is in line with the notion that time-reversal symmetry breaking can reduce entropy production Benenti et al. (2017). The device’s performance as a refrigerator can also be optimised, but the optimal value of $\sigma$, as in the previous section, depends of the operating voltage and temperature difference. Finally, despite the arguments for the analysis with normalised conductance, it should also be noted that increasing $\xi$ increases overall conductance and therefore thermal machine performance.
@@ V Noise and constancy
The final significant characteristic of a thermal machine that we have not considered so far is noise, or the constancy of power delivery or extraction. The trade-off between power, efficiency, and constancy of a general heat engine has been extensively studied Shiraishi et al. (2016); Pietzonka and Seifert (2018), however the results are largely based on thermodynamic uncertainty relations (TURs)Barato and Seifert (2015); Horowitz and Gingrich (2020); Seifert (2012), which are called into question in the case of detailed balance breaking Macieszczak et al. (2018); Taddei and Fazio (2023). Kinetic uncertainty relations (KURs) Landi et al. (2024) based on dynamical analysis hold under detailed balance breaking, providing an independent bound for noise. However, the KUR bound is notoriously weak close to equilibrium Palmqvist et al. (2025), and therefore not restrictive in the linear response regime.
We thus take a fully dynamical, instead of thermodynamic, approach to the analysis of noise. A general expression for current fluctuations, $\chi$, as a function of four tunnel rates for a SET has the form Landi et al. (2024):
@@
where, for brevity, we define: $\Gamma_{L/R}=\Gamma_{TL/R}+\Gamma_{FL/R}$.
The KUR gives:
$\frac{\chi}{I^{2}}\geq\frac{1}{K}$ (22)
where $K$ is activity, equal to $\Gamma_{T}+\Gamma_{F}$. Combined with the explicit expression for current through the four rates (Eq.46), this gives a bound:
$\frac{2\Gamma_{TL}\Gamma_{FL}\Gamma_{R}^{2}+2\Gamma_{TR}\Gamma_{FR}\Gamma_{L}^{2}+4\Gamma_{TL}\Gamma_{FL}\Gamma_{TR}\Gamma_{FR}+(\Gamma_{L}+\Gamma_{R})^{2}(\Gamma_{TL}\Gamma_{FR}+\Gamma_{TR}\Gamma_{FL})}{\left(\Gamma_{TL}\Gamma_{FR}-\Gamma_{FL}\Gamma_{TR}\right)^{2}}\geq 1$ (23)
The fluctuations of power output are then equal to $\varepsilon\chi$. However, fluctuations of current in the output circuit are detrimental for thermal machine operation even if they are associated with little power (around $\varepsilon=0$), so for our analysis, we will consider the fluctuations of current directly.
Unlike the power and efficiency, both based on mean current through the device, in which all three asymmetry parameters: $\Delta S$, $\gamma$, and $\alpha$, came in in a single quantity, noise, or current fluctuations, has a more direct dependence on the microscopic dynamics of the system, and the three asymmetry parameters, $\Delta S$, $\gamma$, and $\alpha$ all affect it differently. Figure 6a shows the effect of all three asymmetry parameters on the power vs. noise diagrams for a heat engine for a normalised conductance peak height.
@@
A non-zero entropy difference between the charge states, $\Delta S$ splits the dependence for a non-degenerate level into two branches – one with higher power and one with lower for the same noise levels, corresponding to the asymmetric peaks of the thermocurrent (Fig.6a(i)). The tunnel coupling asymmetry coefficient, $\gamma$, affects the power vs. noise relation is a much less profound way (Fig.6(ii)), as, in the absence of detailed balance breaking, it doesn’t affect conductance, and thus output power, but it can be observed that absolute large values of $\gamma$ reduce current fluctuations to a small extent. Finally, the introduction of detailed balance breaking, described by non-zero values of $\alpha$, leads to a dramatic increase of current fluctuations, even in the absence of concurrent tunnel coupling asymmetry and thus the same power output (Fig.6a(iii)). This is in line with the effect of noise amplification with detailed balance breaking, well-known in stochastic thermodynamics Weiss (2003, 2007); Gnesotto et al. (2018). A non-zero value of $\gamma$ would lead to an advantage in power, similarly to an increase of entropy difference.